OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

Part 1: There are no solutions Part 2: y=1 and x=12

OpenStudy (anonymous):

Part 2: (-1)(x-2y=14)=-x+2y=-14 2y=-14 3y=9 5y=-5 y=-1 x+ (-2)(-1) =14 x+2=14 x=12

OpenStudy (anonymous):

y = -1 not 1 sorry on the first post

OpenStudy (radar):

Part 1: I would know the solution was not a valid solution, that the system was inconsistent. x=x x=14+2y x=9-3y since x=x we know that 14+2y=9-3y 5y=-5 y=-1 x=12